# A phenomenological gravitational waveform model of binary black holes incorporating horizon fluxes

### v2

@@ IV Hybrid construction
For constructing a hybrid, the NR waveform has to be ‘stitched’ with the analytical inspiral waveform in a frequency region where both the waveforms have sufficient accuracy, which also demands that the morphology of these waveforms do not deviate significantly from one another. For PN waveforms, one conventionally chooses ISCO of the binary to be the endmost point of validity. The corresponding GW frequency $f_{\rm ISCO}$ is considered to be the highest frequency for PN expansions. In our work, we ensured that the frequency at the stitching region is below this threshold for all the hybrids - which also ensures that the NR calibrations in SEOBNRv2 are minimally present before the stitching starts.
@@
  • A robust hybrid should depend weakly on the small changes to the stitching region. A monochromatic signal would be completely degenerate under a shift in the coalescence time $t_{c}$ and the coalescence phase $\phi_{c}$; this degeneracy is broken by the increase in frequency with time. The start and end of the stitching interval should reflect enough change in the frequency to break this degeneracy.
The first issue is addressed in our work by choosing the stitching region beyond the specified relaxation time for each NR simulation in the SXS catalog, the time interval (in units of $M$) by which the junk radiation dies out. To address the second requirement, MacDonald et al. [70] have recommended the stitching interval to satisfy $\delta\omega/\omega_{m}\geqslant 0.1$, where $\omega_{m}$ is the GW frequency at the midpoint of the stitching interval, and $\delta\omega$ is the change in frequency over the interval. This choice ensures that the residual oscillations in $t_{c}$ with changing $\omega_{m}$ are below $1M$. In our work, we perform the stitching over 4 GW cycles, and we place the stitching region in a way to satisfy this condition. In the literature, the construction of hybrids has been performed both in time  [20, 24, 23] and frequency domains [21, 71]. In our work, we choose the former, primarily because the NR waveforms are available in the time domain, and typically they have too few numbers of GW cycles for a reliable Fourier transform.
@@
    SEOBNRv2 is a time-domain waveform, and the phase and amplitude corrections for TH are calculated in the frequency domain. So, we have to first convert the time-domain data to the Fourier domain to apply the corrections. Since the time-domain data is finite, we use a tapered window function to minimize Gibbs artifacts, called the Planck-taper window [70]:
    The frequency-domain waveform, after correcting for TH, is then converted back to the time domain for hybridization.
@@
    NR data from the SXS catalog sample the GW strain $h^{(\rm NR)}(t)$ in non-uniform timesteps, with a higher sampling rate in regions of higher GW frequency. We resample the data with the coarsest sampling rate by using cubic interpolation.
@@ V.1 Inspiral phase model
The inspiral approximant used in the hybrid data is SEOBNRv2 with corrections in phase and amplitude due to tidal heating, as discussed earlier. Since the final model is a phenomenological one, it requires ready-to-use analytical forms for the phase and amplitude with extra coefficients calibrated to the hybrid data. We write the frequency-domain phase of a waveform in the inspiral as
$\Psi_{\rm\tiny{INS}}(F;\Theta)=$ $\leavevmode\nobreak\ \Psi_{\rm\tiny{TF2}}(F;\Theta)+\Psi_{\rm TH}(F;\Theta)$ (24)
@@
where $\Psi_{\rm\tiny{TF2}}(F;\Theta)$ and $\Psi_{\rm TH}(F;\Theta)$ are 3.5PN phase of TaylorF2 (described in Appendix A) and the TH correction (Eq. (11)), respectively. $\sigma_{i}\>\>(i=0-4)$ are 5 phenomenological pseudo-PN parameters. The analytical ansatz in Eq. (24) is the same as in PhenomD inspiral model. Since we want to leverage the modular nature of PhenomD to concatenate its merger-ringdown model with our own inspiral model, the end of the inspiral phase in the current work needs to remain unchanged, which is $F=0.018$. We also note here that one can, in principle, directly add the correction $\Psi_{\rm TH}(F;\Theta)$ to the PhenomD inspiral phase (as is done for the amplitude, discussed in Sec. V.2), without any recalibration. However, since we intend to avoid adding the correction to the NR information present in the last few cycles of inspiral, recalibrating the phase allows a smooth transition from PN to NR. To find the parameter values corresponding to each hybrid waveform, we fit the first derivative of the phase with respect to $F$:
$\Psi^{\prime}_{\rm\tiny INS}(F;\Theta)=$ $\leavevmode\nobreak\ \Psi^{\prime}_{\rm\tiny{TF2}}(F;\Theta)+\Psi^{\prime}_{\rm TH}(F;\Theta)$ (25)
@@ V.2 Inspiral and intermediate amplitude model
  • Solving the aforementioned set of equations, one finds the form of $\tilde{A}_{\rm int}$ which smoothly connects the intermediate region with inspiral and merger.
@@
In Appendix B, we describe the evaluation of the peak frequency $f_{\rm peak}$ and the ringdown frequency $f_{\rm RD}$ as functions of the BH masses and spins.
@@ V.3 Correspondence between the phenomenological and physical parameters
Figure 3: Phase, phase derivative, and phase difference between the hybrids (same as in Fig. 1), individual fits obtained from Eq. (25), and the final model prediction. Top panels: Inspiral phase $\Psi(Mf)$. Middle panels: $-\Psi^{\prime}(Mf)=-\partial\Psi/\partial(Mf)$ plotted against $Mf$. Bottom panels: $\Delta\Psi_{\rm Fit}=(\Psi_{\rm Hybrid}-\Psi_{\rm Fit})$, and $\Delta\Psi_{\rm Model}=(\Psi_{\rm Hybrid}-\Psi_{\rm Model})$. All the plots are shown within the inspiral frequency range, $0.0035\leqslant Mf\leqslant 0.018$.
@@
For each $\sigma_{i}$, this polynomial generates a set $\{\lambda^{i}_{jk}\}$ to build the correspondence between physical and phenomenological parameter sets. Figure 2 shows the surfaces of $\sigma_{i}$ as functions of $\{\eta,\chi_{\rm PN}\}$ with the ansatz in Eq. (30). These best-fit surfaces generate a constant set of $\{\lambda^{i}_{jk}\}$ which define the inspiral model, along with the TaylorF2 phase and the TH correction. In Appendix C, we list the coefficient values corresponding to each $\sigma_{i}$. A schematic flowchart for generating the inspiral phase model starting from the intrinsic binary parameters can be described as:
@@ VI Faithfulness of the model
Figure 6: Mismatches (%) between PhenomD_Horizon and PhenomD. (a) Mismatches with only the phase correction but no amplitude correction in PhenomD_Horizon. (b) Mismatches with only the amplitude correction without any phase correction. (c) Mismatches with corrections in both the phase and amplitude. The plots show that the dominant contributor to these mismatches is the phase correction.
@@ VII Model comparison
Figure 6 shows the mismatches between PhenomD_Horizon and PhenomD, keeping $\chi_{1}=\chi_{2}$. We explore the mismatches within the parameter space $q\in[1,8]$ and $\chi_{1},\chi_{2}\in[-0.9,0.9]$, for six values of the binary mass, $M/M_{\odot}\in[10,20,30,40,50,60]$. In Fig. 6, we show the mismatches including only the phase correction - i.e. using the recalibrated inspiral phase, but keeping the IMR amplitude the same as PhenomD. Since the phase correction due to TH written in Eq. (11) is $\propto 1/\eta$ at the leading order, higher mismatches are found for higher mass ratios, for a fixed value of the spins. With increasing spins, mismatches rise due to the increased significance of TH. The mismatch reaches a maximum of $\sim 16\%$ for $30M_{\odot}$ binaries with $q\gtrsim 7$ and $\chi_{1}=\chi_{2}\gtrsim 0.8$, and a minimum of $\sim 10^{-5}\%$ for heavier binaries. For $M\leqslant 30M_{\odot}$, we see substantial regions of the parameter space with mismatches between 1-10% for spin magnitudes of $|\chi_{1}|=|\chi_{2}|\gtrsim 0.5$. For $M>30M_{\odot}$, the 1% level is crossed only for $q\gtrsim 4$.
@@
In Fig. 6, we use the complete IMRPhenomD_Horizon model, by including both the phase and amplitude corrections. The contours of mismatches $\geqslant 0.1\%$ are almost the same as in Fig. 6, but amplitude corrections elevate them above 0.01% throughout other regions of the parameter space. However, mismatches $<0.1\%$ are below the accuracy level of the model, for which we considered a 1% tolerance level of modeling errors (Fig. 5). This suggests that including the amplitude correction in the TH corrected model has negligible effect compared to the phase correction in GW data analysis.
To demonstrate the improvement in accuracy that our model introduces over PhenomD, we use 219 NR waveforms from the SXS catalog and compare them against PhenomD and PhenomD_Horizon separately by computing mismatches. In Fig. 7, we show the distributions of the mismatches, by using the Advanced LIGO ZDHP noise curve. We plot histograms of $\log_{10}(1-\mathcal{M})$, where $(1-\mathcal{M})$ is the average mismatch in the binary mass range $12M_{\odot}\leqslant M\leqslant 100M_{\odot}$. The dashed lines show the medians of the two histogram plots, where a slight overall improvement is seen. The fractional shift in the medians (denoted by an overbar), is
@@
showing an improvement of $\sim 4\%$. Figure 8 shows similar histograms with a flat noise curve. In this case, also we see a shift in the median value towards lower mismatches, with an improvement of $\sim 2\%$.
@@ VIII Discussion and future work
We have presented a phenomenological gravitational waveform model for binary black holes that carries the signatures of tidal heating of the holes explicitly throughout the binary evolution, starting from the early inspiral phase and ending at their merger. In the frequency domain, it contains corrections due to this effect both in the phase and amplitude. The entire IMR model for the phase is divided into two parts. The inspiral phase model, defined by the frequency range $0.0035\leqslant Mf\leqslant 0.018$, is a recalibrated version of IMRPhenomD [24, 25] to account for the modifications due to tidal heating. The merger-ringdown parts, defined as $Mf\geqslant 0.018$ in terms of the dimensionless frequency, is the same as PhenomD since this region contains purely NR data, which inherently incorporates the dephasing due to tidal heating. The amplitude corrections are added by dividing the model into three parts. The inspiral model, with $0.0035\leqslant Mf\leqslant 0.014$, is constructed by directly adding the correction term to the amplitude model of PhenomD in this region. The intermediate region, $0.014\leqslant Mf\leqslant Mf_{\rm peak}$, $f_{\rm peak}$ being the frequency corresponding to the peak amplitude, is approximated as a polynomial in $f$ and evaluated by imposing $C^{(1)}$ continuity at the two ends and a collocation point at their mid-frequency. The amplitude model is identical to PhenomD for $f>f_{\rm peak}$. The final model is calibrated within the range in the mass ratio $q\in[1,8]$ and the effective spin parameter $\chi_{\rm PN}\in[-0.79,0.72]$.
@@
We conducted a study of faithfulness, by calculating mismatches between the model and a validation dataset of 40 hybrid waveforms constructed with tidal heating included. We demonstrated that modeling errors do not lead to mismatches higher than 1% (barring a single outlier), with most of the mismatch values lying around 0.1%. This standard of faithfulness is comparable to PhenomD, which also considered a maximum tolerance level of 1% mismatch.
@@
How the new model differs from its parent model PhenomD was assessed by calculating mismatches between them within the calibration range of the intrinsic parameter space, for binary masses ranging from $10M_{\odot}$ to $60M_{\odot}$. Phase corrections in the inspiral result in mismatches between these two models as high as $\sim 16\%$, whereas amplitude corrections alone are unable to produce mismatches above $\sim 0.001\%$. With increasing mass asymmetry, the impact of the phase correction increases, but the amplitude correction becomes less effective. Combined with the fact that the former has more significance in GW searches and PE, this reciprocal nature of these two correction terms leaves the latter mostly superfluous for practical purposes.
@@
We compared the model with a set of 219 NR waveforms from the SXS catalog [54], and show the mismatch distribution along with the same for PhenomD, averaged over the binary mass range $12M_{\odot}\leqslant M\leqslant 100M_{\odot}$. With the Advanced LIGO ZDHP noise curve, we see an improvement of $\sim 4\%$ in the medians of the two distributions. With a flat noise curve, the improvement is relatively modest, about 2%. These overall improvements point towards a mild yet non-negligible increase in waveform accuracy.
@@
While this work has been in progress, Ref. [43] has calculated the horizon flux expressions with more accurate terms up to the relative 1.5PN order, which agree with the perturbative calculations in the extreme mass-ratio limit [78]. We note that the expressions in Ref. [43] should be used in more accurate models that are valid up to high mass ratios.

### v3

@@ IV Hybrid construction
For constructing a hybrid, the NR waveform has to be ‘stitched’ with the analytical inspiral waveform in a frequency region where both the waveforms have sufficient accuracy, which also demands that the morphology of these waveforms do not deviate significantly from one another. For PN waveforms, one conventionally chooses the innermost circular orbit (ISCO) of the binary to be the endmost point of validity. The corresponding GW frequency $f_{\rm ISCO}$ is considered to be the highest frequency for PN expansions. In our work, we ensured that the frequency at the stitching region is below this threshold for all the hybrids.
@@
  • A robust hybrid should depend weakly on small changes to the stitching region. A monochromatic signal would be completely degenerate under a shift in the coalescence time $t_{c}$ and the coalescence phase $\phi_{c}$; this degeneracy is broken by the increase in frequency with time. The start and end of the stitching interval should reflect enough change in the frequency to break this degeneracy.
The first issue is addressed in our work by choosing the stitching region beyond the specified relaxation time for each NR simulation in the SXS catalog, the time interval (in units of $M$) by which the junk radiation dies out. To address the second requirement, MacDonald et al. [63] have recommended the stitching interval to satisfy $\delta\omega/\omega_{m}\gtrsim 0.05$, where $\omega_{m}$ is the GW frequency at the midpoint of the stitching interval, and $\delta\omega$ is the change in frequency over the interval. In our work, we perform the stitching over 4 GW cycles, and we place the stitching region in a way to satisfy this condition. In the literature, the construction of hybrids has been performed both in time  [20, 24, 23] and frequency domains [21, 78]. In our work, we choose the former, primarily because the NR waveforms are available in the time domain, and typically they have too few GW cycles for a reliable Fourier transform.
@@
    NR data from the SXS catalog sample the GW strain $h^{(\rm NR)}(t)$ in non-uniform timesteps, with a higher sampling rate in regions of higher GW frequency. We resample the data with the coarsest sampling rate by using cubic interpolation. Since our modeling only involves the inspiral parts of the waveforms, this choice does not lead to any loss of information.
@@
    SEOBv2_Horizon is generated with the same set of intrinsic parameters that forms the inspiral part of the hybrids, with the same timesteps.
@@ V.1 Inspiral phase model
The inspiral approximant used in the hybrid data is SEOBv2_Horizon, as discussed earlier. Since the final model is a phenomenological one, it requires ready-to-use analytical forms for the phase and amplitude with extra coefficients calibrated to the hybrid data. We write the frequency-domain phase of a waveform in the inspiral as
$\Psi_{\rm INS}(F;\Theta)=$ $\leavevmode\nobreak\ \Psi_{\rm{TF2}}(F;\Theta)+\Psi_{\rm TH}(F;\Theta)$ (22)
@@
where $\Psi_{\rm{TF2}}(F;\Theta)$ and $\Psi_{\rm TH}(F;\Theta)$ are 3.5PN phase of TaylorF2 (described in Appendix A) and the TH correction (Eq. (10)), respectively. $\sigma_{i}\>\>(i=0-4)$ are 5 phenomenological pseudo-PN parameters. The analytical ansatz in Eq. (22) is the same as in PhenomD inspiral model. Since we want to leverage the modular nature of PhenomD to concatenate its merger-ringdown model with our own inspiral model, the end of the inspiral phase in the current work needs to remain unchanged, which is $F=0.018$. We also note here that one can, in principle, directly add the correction $\Psi_{\rm TH}(F;\Theta)$ to the PhenomD inspiral phase (as is done for the amplitude, discussed in Sec. V.2), without any recalibration. However, since we intend to avoid adding the correction to the NR information present in the last few cycles of inspiral, recalibrating the phase allows a smooth transition from PN to NR. To find the parameter values corresponding to each hybrid waveform, we fit the first derivative of the phase with respect to $F$:
$\Psi^{\prime}_{\rm INS}(F;\Theta)=$ $\leavevmode\nobreak\ \Psi^{\prime}_{\rm{TF2}}(F;\Theta)+\Psi^{\prime}_{\rm TH}(F;\Theta)$ (23)
@@ V.2 Inspiral and intermediate amplitude model
  • Solving the aforementioned set of equations, one finds the form of $\tilde{A}_{\rm int}$, which smoothly connects the intermediate region with inspiral and merger.
@@
In Appendix C, we describe the evaluation of the peak frequency $f_{\rm peak}$ and the ringdown frequency $f_{\rm RD}$ as functions of the BH masses and spins.
@@ V.3 Correspondence between the phenomenological and physical parameters
Figure 5: Phase, phase derivative, and phase difference between the hybrids (same as in Fig. 3), individual fits obtained from Eq. (23), and the final model prediction. Top panels: Inspiral phase $\Psi(Mf)$. Middle panels: $\Psi^{\prime}(Mf)=\partial\Psi/\partial(Mf)$ plotted against $Mf$. Bottom panels: $\Delta\Psi_{\rm Fit}=(\Psi_{\rm Hybrid}-\Psi_{\rm Fit})$, and $\Delta\Psi_{\rm Model}=(\Psi_{\rm Hybrid}-\Psi_{\rm Model})$. All the plots are shown within the inspiral frequency range, $0.0035\leqslant Mf\leqslant 0.018$.
@@
For each $\sigma_{i}$, this polynomial generates a set $\{\lambda^{i}_{jk}\}$ to build the correspondence between physical and phenomenological parameter sets. Figure 4 shows the surfaces of $\sigma_{i}$ as functions of $\{\eta,\chi_{\rm PN}\}$ with the ansatz in Eq. (28). These best-fit surfaces generate a constant set of $\{\lambda^{i}_{jk}\}$ which define the inspiral model, along with the TaylorF2 phase and the TH correction. In Appendix D, we list the coefficient values corresponding to each $\sigma_{i}$. A schematic flowchart for generating the inspiral phase model starting from the intrinsic binary parameters can be described as:
@@ 20 SXS:BBH:0298 7 0 0
[…]
@@ VII Model comparison
First, to look at how the TH corrections impact an analytical waveform model, we compare 3.5PN TaylorF2, with the frequency-domain phase defined in Appendix A, with and without the TH corrections in the phase and amplitude. Figure 8 shows the corresponding mismatches, keeping $\chi_{1}=\chi_{2}$. We explore the mismatches within the parameter space $q\in[1,8]$ and $\chi_{1},\chi_{2}\in[-0.9,0.9]$, for six values of the binary mass, $M/M_{\odot}\in[10,20,30,40,50,60]$. The frequency range for the mismatch calculation is $f\in[10,0.018/M]$ (Hz). Since the phase correction due to TH written in Eq. (10) is $\propto 1/\eta$ at the leading order, higher mismatches are found for higher mass ratios, for a fixed value of the spins. With increasing spins, mismatches rise due to the increased significance of TH. Figure 8 is similar to Fig. 8 in every way only except the upper cutoff frequency, which is chosen to be the ISCO of the corresponding merger remnant KBH [82]. Before we compare PhenomD_Horizon and PhenomD, we note that there are some differences in modeling and calibrating the two waveforms apart from just the addition of the correction due to horizon fluxes. First, PhenomD was calibrated up to $q=18$, with a different set of hybrid waveforms. Second, the quadratic and cubic spin terms were absent in the TaylorF2 phasing, which we take into account at the 3PN and 3.5PN orders, respectively. Thus, to discern the physical difference that PhenomD_Horizon introduces, we create another model (IMRPhenomD_Recal) by leaving out the TH corrections in the phase and amplitude, but keeping every other aspect unchanged. We expect this model, being free from any modeling differences with PhenomD_Horizon, to produce mismatches comparable to Fig. 8. We plot them in Fig. 8 and, indeed, we see that the mismatches exhibit a qualitative similarity. However, since the upper cutoff is fixed at the inspiral cutoff of PhenomD ($Mf=0.018$), and the hybrids used for calibration of both PhenomD_Horizon and PhenomD_Recal are stitched to NR data before that, we see some differences between these two plots. Specifically, in Fig. 8 we show the maxima, minima, and the medians of mismatches for each of the subplots in Figs. 88, and 8. Overall, the analytical waveforms show higher mismatches than the phenomenological ones, which can be ascribed to the presence of NR data in the hybrids between the stitching frequency and the inspiral upper cutoff of $Mf=0.018$. With increasing binary mass, the differences are higher, which is expected since the similarity between PhenomD_Horizon and PhenomD_Recal are more towards higher frequencies.
Figure 9 shows the mismatches between IMRPhenomD_Horizon and the original IMRPhenomD  [^3] This should be distinguished from PhenomD_Recal that we used for Fig. 8. We constructed PhenomD_Recal to find out how much mismatch arises solely due to TH. In Fig. 9, however, the mismatch values with PhenomD originate both from modeling difference as well as TH. , now in the frequency range $f\in[10,0.3/M]$ (Hz). In Fig. 9, we show the mismatches including only the phase correction - i.e. using the recalibrated inspiral phase, but keeping the IMR amplitude the same as PhenomD. The mismatch reaches a maximum of $\sim 14\%$ for $30M_{\odot}$ binaries with $q\gtrsim 7$ and $\chi_{1}=\chi_{2}\gtrsim 0.8$, and a minimum of $\sim 10^{-5}\%$ for heavier binaries. We see substantial regions of the parameter space with mismatches between 1-10% for spin magnitudes of $|\chi_{1}|=|\chi_{2}|\gtrsim 0.5$, and $q\gtrsim 4$.
@@
In Fig. 9, we use the complete IMRPhenomD_Horizon model, by including both the phase and amplitude corrections. The distribution of mismatches in this figure shows great similarity with Fig. 9. In Fig. 9, we plot the maxima, minima, and medians of mismatches for all the subfigures, the same way we did for Fig. 8. This suggests that including the amplitude correction in the TH corrected model has negligible effect compared to the phase correction in GW data analysis.
To demonstrate the improvement in accuracy that our model introduces over PhenomD, we use 219 NR waveforms from the SXS catalog and compare them against PhenomD and PhenomD_Horizon separately by computing mismatches. In Fig. 10, we show the distributions of the mismatches by using the Advanced LIGO ZDHP noise curve. We plot histograms of $\log_{10}(1-\mathcal{M})$, where $(1-\mathcal{M})$ is the average mismatch in the binary mass range $12M_{\odot}\leqslant M\leqslant 100M_{\odot}$. The dashed lines show the medians of the two histogram plots, where a slight overall improvement is seen. The fractional shift in the medians (denoted by an overbar), is
@@
showing an improvement of $\sim 4\%$. Figure 11 shows similar histograms with a flat noise curve. In this case, we also see a shift in the median value towards lower mismatches, with an improvement of $\sim 1.5\%$.
@@ VIII Discussion and future work
The main objective of this work is to present a phenomenological gravitational waveform model for binary black holes that carries the signatures of tidal heating of the holes explicitly throughout the binary evolution, starting from the early inspiral phase and ending at their merger. We also demonstrate, in Sec VI, that inclusion of corrections due to tidal heating of black holes, in the phase and amplitude, rewards an overall improvement in agreement between the point-particle analytical description (here it is uncalibrated SEOBNRv2) of a compact binary coalescence and numerical relativity data, especially for high mass-ratios and high spins. In the frequency domain, the waveform model presented here contains corrections due to this effect both in the phase and amplitude. The entire IMR model for the phase is divided into two parts. The inspiral phase model, within the frequency range $0.0035\leqslant Mf\leqslant 0.018$, is a recalibrated version of IMRPhenomD [24, 25] to account for the modifications due to tidal heating. The merger-ringdown parts, defined as $Mf\geqslant 0.018$ in terms of the dimensionless frequency, are the same as PhenomD since this region contains purely NR data, which inherently incorporates the dephasing due to tidal heating. The amplitude corrections are added by dividing the model into three parts. The inspiral model, with $0.0035\leqslant Mf\leqslant 0.014$, is constructed by directly adding the correction term to the amplitude model of PhenomD in this region. The intermediate region, $0.014\leqslant Mf\leqslant Mf_{\rm peak}$, $f_{\rm peak}$ being the frequency corresponding to the peak amplitude, is approximated as a polynomial in $f$ and evaluated by imposing $C^{(1)}$ continuity at the two ends and a collocation point at their mid-frequency. The amplitude model is identical to PhenomD for $f>f_{\rm peak}$. The final model is calibrated within the range in the mass ratio $q\in[1,8]$ and the effective spin parameter $\chi_{\rm PN}\in[-0.79,0.72]$.
@@
We conducted a study of faithfulness by calculating mismatches between the model and a validation dataset of 40 hybrid waveforms constructed with tidal heating included. We demonstrated that modeling errors do not lead to mismatches higher than 1% (barring a single outlier), with most of the mismatch values lying around 0.1%. This standard of faithfulness is comparable to PhenomD, which also considered a maximum tolerance level of 1% mismatch.
@@
How the new model differs from its parent model PhenomD was assessed by calculating mismatches between them within the calibration range of the intrinsic parameter space, for binary masses ranging from $10M_{\odot}$ to $60M_{\odot}$. Phase corrections in the inspiral result in mismatches between these two models as high as $\sim 14\%$, whereas amplitude corrections alone are unable to produce mismatches above $\sim 0.001\%$. With increasing mass asymmetry, the impact of the phase correction increases, but the amplitude correction becomes less effective. Combined with the fact that the former has more significance in GW searches and PE, this reciprocal nature of these two correction terms leaves the latter mostly superfluous for practical purposes.
@@
We compared the model with a set of 219 NR waveforms from the SXS catalog [57], and show the mismatch distribution along with the same for PhenomD, averaged over the binary mass range $12M_{\odot}\leqslant M\leqslant 100M_{\odot}$. With the Advanced LIGO ZDHP noise curve, we see an improvement of $\sim 4\%$ in the medians of the two distributions. With a flat noise curve, the improvement is relatively modest, about 1.5%. These overall improvements point towards a mild yet non-negligible increase in waveform accuracy.